# Exploring Prompting Methods for Mitigating Class Imbalance through Synthetic Data Generation with Large Language Models

### v2

@@ Abstract
Large language models (LLMs) have demonstrated impressive in-context learning capabilities across various domains. Inspired by this, our study explores the effectiveness of LLMs in generating realistic tabular data to mitigate class imbalance. We investigate and identify key prompt design elements such as data format, class presentation, and variable mapping to optimize the generation performance. Our findings indicate that using CSV format, balancing classes, and employing unique variable mapping produces realistic and reliable data, significantly enhancing machine learning performance for minor classes in imbalanced datasets. Additionally, these approaches improve the stability and efficiency of LLM data generation. We validate our approach using six real-world datasets and a toy dataset, achieving state-of-the-art performance in classification tasks. The code is available here.
Figure 1: Generation results on an imbalanced toy dataset. The major class is shown in blue, and the minor class in red. Our approach via in-context learning of LLM achieves (1) clear class boundaries, (2) accurate feature correlations, (3) proper value ranges, (4) well-represented relationships between categorical and numerical variables (last row), and (5) better diversity compared to the fine-tuned GReaT [4] model. Detailed explanations and complete results are provided in Appendix E.
@@ 1 Introduction
Tabular data plays a crucial role in various classification tasks, such as disease diagnosis [1] and financial analysis [3, 24]. However, challenges such as data scarcity and class imbalance, particularly for rare yet crucial events, can severely degrade the performance of machine learning (ML) models. To address these, synthetic data generation methods have been developed to augment tabular datasets cost-effectively. Traditional methods such as SMOTE and its variants [6, 12, 20] focus on generating data for minority classes to alleviate class imbalance. Recently, advanced deep learning-based generation models have shown promising results, such as TVAE [30], CTAB-GAN [33], and TabDDPM [18].
@@
Large language models (LLMs) have also demonstrated significant potential in generating realistic tabular data, showing robust performance across numerical and categorical data [4, 31]. However, these models often require extensive, data-specific fine-tuning, which can lead to overfitting on major classes or dominant feature values when applied to small and imbalanced tabular datasets. An alternative approach leverages the in-context learning capabilities of LLMs. Extensive studies have demonstrated that LLMs can solve complex problems, learn and mimic patterns from input data, and generate new data without parameter updates [13, 17, 19, 28, 29, 32]. This capability holds significant potential for generating realistic and reliable tabular data to address class imbalance.
However, the in-context learning capabilities of LLMs largely depend on a prompt design [10, 26, 28, 32], while these methods are not directly applicable to tabular data generation, especially for imbalanced class scenarios. This is because tabular data is not inherently in plain text and involves unique challenges such as identifying feature correlations and accurately representing underrepresented attributes. Therefore, carefully crafted prompt methods specifically for imbalanced tabular data are essential to maximize LLM capabilities in generating realistic data effectively. While several studies have highlighted the potential of LLMs for generating synthetic tabular data using in-context learning [25, 31], they have not sufficiently explored various prompt design elements, such as data format or class presentation. This study aims to investigate the effectiveness of LLMs in resolving class imbalance problems and identify the optimal prompts to maximize this effectiveness.
Figure 2: Comparison of ML classification performance when adding synthetic data to the Travel dataset. The results are averaged across four classifiers, XGBoost, CatBoost, LightGBM, and Gradient Boosting Classifier. Each classifier is run five times, resulting in a total of 20 experiments.
To answer these questions, we explore various prompt methods and analyze the generated results to identify specific prompt structures and templates that significantly improve the quality of synthetic tabular data generated by LLMs, thereby enhancing ML classification performance. We propose several prompt design elements, including data format, class presentation methods, variable mapping, and task specification. Our findings indicate that providing data in CSV format, balancing and grouping classes, and applying a unique variable mapper to distinguish different variables significantly improve the quality and efficiency of synthetic data generation, underscoring their effectiveness in optimizing LLM performance. As shown in Fig. 1, the data generated using this prompt design exhibit the most distinct boundaries between target classes, accurate feature correlation, and variable ranges.
Moreover, this approach significantly improves the classification performance of ML models when applied to real-world datasets. As shown in Fig. 2, baseline methods exhibit low sensitivity, indicating that they fail to accurately generate minority class samples due to inherent class biases. In contrast, our method overcomes these, achieving high sensitivity and balanced performance for all classes.
In summary, our contributions are:
@@
  • Our study explores and analyzes the effectiveness of LLMs in generating realistic synthetic tabular data via in-context learning, offering guidelines for designing prompts tailored to effectively exploit LLMs for this task, particularly addressing class imbalance.
  • We propose simple yet effective prompting methods that generate high-quality class-specific tabular data, maintaining feature correlation and value ranges.
  • The proposed approach significantly improves ML classification performance and enhances data generation efficiency and stability via in-context learning.
  • The proposed approach is generally applicable and easy to implement for any tabular classification task with minimal pre-processing, providing publicly available code.
@@
  • We validate the effectiveness of our approach through extensive evaluation across six real-world public tabular data classification datasets and a toy dataset.
@@ 2 Method
In this study, we explore various prompt design options to maximize the in-context learning capability of LLMs for generating high-quality tabular data. Our goal is to propose an optimal approach that reliably and efficiently produces realistic tabular data to enhance ML classification performance, especially in the context of class imbalance. We start with a tabular dataset $T\in\mathds{R}^{n\times m}$, comprising $n$ instances, each with $m$ variables. We aim to generate a synthetic dataset, $\hat{T}\in\mathds{R}^{n^{\prime}\times m}$, which accurately mimics the characteristics of $T$, including class conditions, feature correlations, and value ranges. To achieve this, we explore several key prompt design elements: Data format, class presentation methods, variable mapping, and task specification.
We evaluate these design choices based on the quality of the generated data, with a primary focus on ML classification performance and generation efficiency, as detailed in Section 3. Our carefully crafted prompt, as depicted in Fig.3, begins with optional variable descriptions followed by multiple sets of data samples. Each set consists of feature names and samples grouped by target class. By adding feature names at the end of the prompt as a trigger, the LLM generates realistic data that improves ML performance and maintains feature correlations similar to the original data. Prompt examples and an illustration of the ablated versions are provided in Appendices I.5 and I.1, respectively.
@@ 2.1 Data format: Sentence vs CSV style
We investigate effective data formatting methods to represent tabular data within prompts, ensuring that LLMs can interpret the demonstrations and generate new data accurately. Tabular data can be formatted as plain text or as comma-separated values (CSV style). Let the feature name of the $k$-th variable be $v_{k}$, and the $i$-th observation of this variable be $o_{k,i}$. According to previous work [4], this value can be transformed into a sentence-style representation as $[v_{k},``is",o_{k,i},``,"]$. In contrast, the CSV style can be expressed as $[o_{k,i},``,"]$, presenting only values, with variable names for all variables specified as $\bm{v}=[v_{1},``,",v_{2},``,",\cdots,v_{m}]$ at the beginning of the data sample presentation. Here, the sentence style redundantly uses variable names and “is" tokens for every value, leading to higher token usage and computational cost than the CSV style. The CSV-style format allows for a higher volume of in-context learning examples within the same token constraints. Thus, our method utilizes a CSV-style format to maximize the number of examples, as providing a large number of data samples is crucial for ensuring sufficient representation of the original dataset in in-context learning.
@@ 2.2 Class presentation
Class presentation methods aim to ensure that all target classes are adequately represented within a prompt, particularly addressing class imbalance for minority classes.
Single-class vs. Multi-class generation    When generating data with class conditions, a primary consideration is whether to generate data for one class at a time or for all classes simultaneously. Generating data for a single class allows the model to focus on specific class characteristics, whereas generating data for multiple classes simultaneously enables comparison and contrast across classes. Our experiments indicate that generating data for multiple classes simultaneously, in a structured manner (explained subsequently), produces samples that more accurately represent each class compared to single-class generation. This suggests that generating data with an awareness of the characteristics of other classes is more effective.
Original vs. Balanced class ratio    When constructing a prompt to generate multiple classes simultaneously, naive random sampling from the original tabular dataset $T$ often results in an imbalanced ratio of samples for each target class. To address this, we consider balancing the sample numbers for minor and major classes rather than adhering to the original data distribution. Our findings indicate that balancing the sample numbers significantly improves the quality of generated data for minor classes, demonstrating that training with balanced data helps LLMs learn the true distributions of target classes more effectively.
Listing vs. Class grouping    Within a prompt, the data examples can either be listed in the order they are sampled without class distinction or grouped by class. Empirical evidence suggests that grouping by class leads to a better representation of class-specific feature correlations in the generated data. This grouping emphasizes contrasts between different classes and similarities within each group. As a result, the LLM generates more distinct samples for each class and more coherent samples within each class. Additionally, group-wise prompting can effectively condition LLMs to generate data with specific values. Without grouping, it is challenging to constrain LLMs to generate samples with specific attributes, often requiring extensive iterations until the desired samples are produced, which is inefficient. However, when examples are provided in groups by class, LLMs learn this and generate samples for each class in groups. Thus, grouping is crucial for conditioning the LLM to generate data with specific values and conditions. Moreover, grouping can be extended to condition on any arbitrary features other than target classes by grouping samples based on specific requirements.
@@
As with other LLM-based approaches [4], our approach allows minimal data preprocessing, maintaining the original feature names and values. This significantly reduces the need for complex data manipulations often required in existing methods [18, 34] and helps to preserve the integrity of the raw data. However, one critical consideration arises when a dataset includes a large number of categorical variables that share the same values (e.g., extensive boolean variables). In this case, the data examples in the prompt can become monotonous and repetitive, making it challenging for the model to distinguish between different variables. This can significantly degrade the quality of the generated data. Moreover, when the LLM receives such input, the number of data samples generated at once can significantly decrease, and the generation efficiency can greatly decline.
To address this, we propose a unique variable mapping method, which remaps the values of each categorical variable to distinct random alphanumeric strings, as depicted in Fig. 3. Uniform substitution is applied across the dataset, maintaining the integrity of the data structure while introducing necessary variation. For example, consider a categorical variable $v_{1}$ with values of ‘t’ and ‘f’. Our approach remaps these values to distinct three-character strings, such as ‘JY0’ and ‘GGN,’ respectively, and applies them consistently across all data points in the dataset. Other categorical variables are similarly remapped to unique values. The transformed dataset is then used as in-context learning demonstrations. Although these new values do not hold inherent semantic meaning, they effectively represent categorical distinctions as symbols, allowing the LLM to identify and utilize patterns within them. This approach ensures that each variable has a unique representation, making the variables clearly distinguishable and contributing to more accurate and efficient data generation.
@@ 2.4 Task specification
A crucial aspect of the prompt design is to clearly define the task within the prompt, guiding the LLM to generate synthetic data as intended. To achieve this, the following options can be considered:
Explicit instruction vs. Completion triggering    We can clarify the task to the LLM by including explicit instructions to “generate new data samples." However, crafting specific instructions can be challenging, as LLMs are sensitive to minor variations in the prompt, making this approach complex for the tabular data generation task. A simpler and more effective method is providing input patterns that LLM can mimic to generate new data, only using data samples without explicit instruction. For example, let the $i$-th set of data samples for all $c$ classes be formatted as $\bm{S}_{i}=[\bm{v},\bm{s}_{i,1},\cdots,\bm{s}_{i,c}]$, where $\bm{s}_{i,j}$ consists of $n$ data samples for the $j$-th class. Here, the feature name $\bm{v}$ works as a header to indicate the start of a set. We input $t$ such sets in a prompt, $[\bm{S}_{1},\bm{S}_{2},\ldots,\bm{S}_{t}]$, creating a pattern and expecting the LLM to generate $\bm{S}_{t\texttt{+}1}=[\bm{v},\bm{s}_{t\texttt{+}1,1},\cdots,\bm{s}_{t\texttt{+}1,c}]$. To ensure this, we add a trigger to induce the LLM to complete the sequence. That is, we include the header of the new set $[\bm{v}]$ at the end of the prompt, which cues the LLM to generate the rest, $[\bm{s}_{t\texttt{+}1,1},\cdots,\bm{s}_{t\texttt{+}1,c}]$, learning the pattern from the input data that these groups are followed by $\bm{v}$. Thus, by creating a pattern in the input prompt, we can condition the LLM to generate a specific number ($c\times n$) of samples for each target class.
Providing contextual information    Optionally, including contextual information in the prompt can help the LLM understand the data structure and generate accurate samples. Therefore, if available, we add line-by-line variable descriptions at the beginning of the prompt. Additionally, one can consider clarifying the target class values at the beginning of each group. However, we observe that treating the target variable as one of the attributes and placing it as the first variable yields better performance, helping the LLM generate more coherent and accurate samples.

### v4

# EPIC: Effective Prompting for Imbalanced-Class Data Synthesis in Tabular Data Classification via Large Language Models

@@ Abstract
Large language models (LLMs) have demonstrated remarkable in-context learning capabilities across diverse applications. In this work, we explore the effectiveness of LLMs for generating realistic synthetic tabular data, identifying key prompt design elements to optimize performance. We introduce EPIC, a novel approach that leverages balanced, grouped data samples and consistent formatting with unique variable mapping to guide LLMs in generating accurate synthetic data across all classes, even for imbalanced datasets. Evaluations on real-world datasets show that EPIC achieves state-of-the-art machine learning classification performance, significantly improving generation efficiency. These findings highlight the effectiveness of EPIC for synthetic tabular data generation, particularly in addressing class imbalance. Our source code for our work is available here.
Figure 1: Generation results on an imbalanced toy dataset with majority and minority classes. Our approach, leveraging in-context learning with LLMs, achieves (1) distinct class boundaries, (2) accurate feature correlations, (3) well-matched value ranges, (4) robust numerical-categorical relationships (last row), and (5) comprehensive data distribution coverage, with improvements over its ablated versions and the fine-tuned GReaT model [4]. Complete results are available in Fig. 9.
Figure 2: Comparison of ML classification performance with synthetic data on the Travel dataset. Results are averaged across four classifiers: XGBoost, CatBoost, LightGBM, and gradient boosting classifier, with each classifier run five times. Our method uses the GPT-3.5-turbo model.
@@ 1 Introduction
Tabular data, consisting of mixed variable types such as numerical and categorical variables, represents a widely applicable and essential data format [30]. It plays a crucial role in enhancing decision-making and efficiency in various real-world applications, such as finance, healthcare, manufacturing, and natural sciences [1, 3, 26]. However, challenges such as data scarcity and class imbalance, particularly for rare yet crucial events, often significantly degrade the performance of machine learning (ML) models, often resulting in reduced accuracy for underrepresented classes. To address these challenges, efficient synthetic data generation methods have been developed to augment tabular datasets. Traditional methods, such as synthetic minority oversampling technique (SMOTE) and its variants [5, 11, 21], focus on generating minority class samples to alleviate class imbalance. Recently, advanced generative models, such as TVAE [35], CTAB-GAN [39], and TabDDPM [19], have shown promising results in producing high-quality synthetic tabular data.
@@
Large language models (LLMs) have also demonstrated significant potential for generating realistic tabular data, effectively handling both numerical and categorical variables [4, 37]. However, these models often require extensive, data-specific fine-tuning, which can increase the risk of overfitting to majority classes or dominant feature values, especially in small and highly imbalanced tabular datasets. An alternative approach involves leveraging the in-context learning capabilities of LLMs. Existing work has shown that LLMs can solve complex reasoning tasks, learn and mimic patterns from input data, and augment textual datasets without parameter updates [13, 20, 34].
However, designing effective prompts to optimize this capability for tabular data generation is challenging, especially for imbalanced datasets. This is because tabular data is not naturally expressed in textual form and involves unique challenges, such as identifying feature correlations and accurately representing underrepresented attributes. Therefore, carefully crafted prompt design tailored to synthetic tabular data generation is essential to fully leverage the capabilities of LLMs. While several studies have employed in-context learning with LLMs for synthetic tabular data generation [27, 37], few have conducted comprehensive investigations into optimal prompt designs that significantly impact data quality and generation efficiency.
In response, this study examines key components of prompt design and identifies an effective method for generating high-quality synthetic tabular data, particularly addressing class imbalance. We introduce a novel approach, EPIC, which leverages the in-context learning capabilities of LLMs to produce synthetic tabular data with balanced class representation. EPIC incorporates prompt design strategies such as CSV style formatting, balanced class grouping, and a unique variable mapper, which together contribute to generating synthetic data that accurately represents class-specific distributions and feature correlations, as illustrated in Fig. 1.
Extensive evaluations across six real-world datasets demonstrate that EPIC significantly improves the data quality and generation efficiency, achieving state-of-the-art performance. As shown in Fig. 2, baselines exhibit low sensitivity, struggling to accurately generate minority class samples due to inherent class imbalances. In contrast, EPIC achieves high sensitivity and balanced performance for all classes, underscoring its robustness and practicality for real-world applications.
In summary, our key contributions are as follows:
@@
  • Our study explores the effectiveness of LLMs in generating realistic synthetic tabular data through in-context learning, providing prompt design guidelines to efficiently generate high-quality data while addressing class imbalance.
  • We propose EPIC, a simple yet effective prompting method that uses balanced, grouped data samples with unique variable mapping to generate tabular data that accurately represents both minority and majority classes, preserving feature correlations and overall data distribution.
  • The proposed approach is model-agnostic, generally applicable to various LLMs, and easy to implement for any tabular data with minimal preprocessing requirements.
@@
  • Extensive experiments on six real-world public tabular datasets and one toy dataset demonstrate the effectiveness of our approach, significantly improving ML classification performance and data generation efficiency.
By addressing class imbalance and enhancing classification outcomes, our work contributes to the advancement of the crucial field of tabular data research, significantly impacting various domains.
@@ 2 Method
In this study, we investigate various prompt design components to maximize the in-context learning capabilities of LLMs for generating high-quality tabular data. Our objective is to develop an optimized approach that reliably and effectively produces realistic tabular data, accurately representing both minority and majority classes to improve ML classification performance, especially addressing class imbalance. Formally, given a tabular dataset $T$ of dimensions $n\times m$, with $n$ samples and $m$ variables, we aim to generate a synthetic dataset $\hat{T}$ of dimensions $n^{\prime}\times m$ that accurately reflects key characteristics of $T$, including class-specific attributes, feature correlations and, data distributions.
To achieve this, we explore the following key prompt design elements: data format (Section 2.1), class presentation methods (Section 2.2), variable mapping (Section 2.3), and task specification (Section 2.4). These design choices are evaluated through extensive analyses on the public datasets from diverse domains, focusing on ML classification performance and generation efficiency. Detailed analysis is provided in Section 3, with Fig. 12 in Appendix C.1 illustrating the ablated versions.
Building on these analyses, we introduce EPIC, a structured prompting method designed to guide LLMs in synthetic data generation while effectively addressing class imbalance, as depicted in Fig. 3. EPIC begins with optional variable descriptions, followed by a series of structured data sample sets, each containing balanced samples for each class. Each set consists of feature names and samples organized by target class. To prompt the LLM to generate a corresponding set of synthetic samples, feature names are appended at the end of the prompt to serve as a trigger, leveraging the pattern recognition capabilities of LLMs. The following sections detail each design component of EPIC.
@@ 2.1 Data format: Sentence vs CSV style
We investigate effective data formatting methods to represent tabular data within prompts, ensuring that LLMs can interpret the demonstrations and generate new data accurately. Tabular data can be formatted as plain text or as comma-separated values (CSV style). Let the feature name of the $k$-th variable be $v_{k}$, and the $i$-th observation of this variable be $o_{k,i}$. According to previous work [4], this value can be transformed into a sentence-style representation as $[v_{k},``is",o_{k,i},``,"]$. In contrast, the CSV style can be expressed as $[o_{k,i},``,"]$, presenting only values, with variable names for all variables specified as $\bm{v}=[v_{1},``,",v_{2},``,",\cdots,v_{m}]$ at the beginning of the data sample presentation. Here, the sentence style redundantly uses variable names and “is" tokens for every value, leading to higher token usage and computational cost than the CSV style. The CSV-style format enables a higher volume of in-context learning examples within the same token constraints. Thus, our method utilizes a CSV-style format to maximize the number of examples, as providing a large number of data samples is crucial for ensuring sufficient representation of the original dataset in in-context learning.
@@ 2.2 Class presentation
Class presentation methods aim to ensure adequate representation of all target classes within a prompt, particularly addressing underrepresented classes in imbalanced datasets. The construction of data samples within the prompt is crucial, as it can either worsen or mitigate existing imbalances. To address this, we explore three prompt design options and propose a class balancing and grouping approach that enables accurate and representative generation of data samples across all target classes.
Single-class vs. Multi-class generation    When generating data under class conditions, a primary consideration is whether to generate data for one class at a time or for all classes simultaneously. Single-class generation allows the model to focus on the unique characteristics of a specific class, while multi-class generation enables direct comparison across classes. Our experiments indicate that generating data for multiple classes simultaneously in a structured manner (discussed in the following sections) yields samples that more accurately capture the distinctive features of each class. This finding suggests that generating data with contextual awareness of other class characteristics enhances the representativeness of synthetic samples.
Original vs. Balanced class ratio    When constructing prompts to generate multiple classes simultaneously, simple random sampling from the original dataset $T$ often results in an imbalanced distribution of samples across target classes. This imbalance may cause the LLM to overfit to the majority class, limiting its ability to learn the characteristics of the minority class. To mitigate this, we employ a balanced sampling approach, equalizing sample numbers for all classes rather than strictly following the original class distribution. Our findings reveal that balancing sample sizes substantially improves the quality of generated data for minority classes, indicating that providing balanced data enables LLMs to learn more effectively and replicate the accurate characteristics of all target classes.
Listing vs. Class grouping    Within a prompt, data samples can either be presented sequentially, in the order they are sampled, or grouped by class. Grouping emphasizes contrasts between different classes and reinforces similarities within each group, enabling the LLM to generate more distinct and coherent samples for each group. Empirical evidence suggests that grouping yields a better representation of class-specific feature correlations in the generated data. Class grouping also contributes to constraining the LLM to generate data with specific attributes more efficiently. Without grouping, guiding the LLM to produce samples with targeted attributes can be challenging, often requiring numerous iterations to achieve the desired outcome. In contrast, grouping samples by specific features facilitates conditioning the LLM to generate data in groups with particular attributes or conditions. Although this work primarily focuses on class conditions, this approach can also be applied to feature grouping, enabling controlled generation based on selected criteria.
@@
Unlike existing methods that often require extensive preprocessing and handling of noisy or missing data [19, 40], our approach minimizes these steps, preserving the integrity of the raw data, including original feature names and values, similar to other LLM-based approaches [4]. However, a critical challenge arises when a dataset contains numerous categorical variables with identical values (e.g., extensive boolean variables). In such cases, data examples in the prompt may become monotonous and repetitive, making it difficult for the model to distinguish between variables, which can significantly degrade the quality of the generated data. Moreover, when faced with such repetitive input, LLMs tend to struggle to generate valid samples, leading to a notable decline in generation efficiency.
To address this, we propose a unique variable mapping method, which remaps the values of each categorical variable to distinct random alphanumeric strings, as illustrated in Fig. 3. Uniform substitution is applied across the dataset, maintaining the integrity of the data structure while introducing necessary variation. For example, consider a categorical variable $v_{1}$ with values of ‘t’ and ‘f’. Our approach remaps these values to distinct three-character strings, such as ‘JY0’ and ‘GGN,’ respectively, and applies them consistently across all data points in the dataset. Other categorical variables are similarly remapped to unique values. The transformed dataset is then used as in-context learning demonstrations. Although these new values do not hold inherent semantic meaning, they effectively represent categorical distinctions as symbols, allowing the LLM to identify and utilize patterns within them. This approach ensures that each variable has a unique representation, making the variables clearly distinguishable and leading to more accurate and efficient data generation.
@@ 2.4 Task specification
A crucial aspect of the prompt design is clearly guiding the LLM to generate synthetic data as intended. To achieve this, we consider two options and ultimately find that prompting the LLM to learn the pattern from a structured, consistent format with optional descriptions for variables is more effective for reliable generation than providing explicit instructions.
Explicit instruction vs. Completion triggering    To clarify the task for the LLM, a prompt might include explicit instructions, such as “generate new data samples." However, crafting specific instructions can be challenging, as LLMs are highly sensitive to subtle prompt variations, making it inefficient to determine an optimal phrasing for tabular data generation. A simpler and more effective approach is to provide patterns within the prompt for the LLM to mimic by formatting data samples without explicit instructions, leveraging its advanced pattern recognition capabilities. Formally, let the $i$-th set of data samples across all $c$ classes be structured as $\bm{S}_{i}=[\bm{v},\bm{s}_{i,1},\cdots,\bm{s}_{i,c}]$, where each $\bm{s}_{i,k}$ contains $n$ samples for the $k$-th class. Here, $\bm{v}$ represents feature names and serves as a header to indicate the start of a set. This approach enables the LLM to recognize the pattern in which each set of data samples is structured and consistently begins with $\bm{v}$. We then input $t$ such sets in a prompt, $[\bm{S}_{1},\bm{S}_{2},\ldots,\bm{S}_{t}]$, to establish a recognizable pattern, expecting the LLM to generate the next set, $\bm{S}_{t\texttt{+}1}=[\bm{v},\bm{s}_{t\texttt{+}1,1},\cdots,\bm{s}_{t\texttt{+}1,c}]$. To reinforce this, we place a trigger at the end of the prompt by including the new header $[\bm{v}]$ for $\bm{S}_{t\texttt{+}1}$, signaling the LLM to complete the sequence with $[\bm{s}_{t\texttt{+}1,1},\cdots,\bm{s}_{t\texttt{+}1,c}]$. By structuring the input prompt in this way, we guide the LLM to generate a consistent number of samples, $n$, for each target class.
Providing contextual information    Optionally, adding contextual information to the prompt can help the LLM understand the characteristics of the original dataset and generate accurate samples. Therefore, when available, we include line-by-line variable descriptions at the beginning of the prompt. Additionally, we treat the target class variable as one of the attributes and position it as the first variable to clarify that each group is organized by the target class.